OpenStudy (anonymous):

Stuck on the question Given a right triangle, one leg is 10 in. and the hypotenuse is 15 in., find the length of the other leg using the Pythagorean theorem. (Use sq rt for the square root symbol).

OpenStudy (anonymous):

10^2= ?

OpenStudy (anonymous):

100

OpenStudy (anonymous):

10^2=100 15^2=225 x= the square root of 125, which is what?

OpenStudy (anonymous):

the Pythagorean theorem is \(a^2+b^2=c^2\) where \(a\) and \(b\) are the legs of the triangles and \(c\) is the hypotenuse, so set up an equation: \(hypotenuse=15\)\[leg 1=10\]\[leg 2= unknown\] \[a^2+b^2=c^2\]\[10^2+b^2=15^2\]\[100+b^2=225\]\[b^2=125\]\[b=\sqrt{125}~in~~~or~~~b\approx11.18034~in\]

OpenStudy (anonymous):

5 sqrt 5

OpenStudy (anonymous):

ya